# SWaT: Statistical Modeling of Video Watch Time through User Behavior Analysis

### v1

@@ Abstract.
The significance of estimating video watch time has been highlighted by the rising importance of (short) video recommendation, which has become a core product of mainstream social media platforms. Modeling video watch time, however, has been challenged by the complexity of user-video interaction, such as different user behavior modes in watching the recommended videos and varying watching probabilities over the video horizon. Despite the importance and challenges, existing literature on modeling video watch time mostly focuses on relatively black-box mechanical enhancement of the classical regression/classification losses, without factoring in user behavior in a principled manner. In this paper, we for the first time take on a user-centric perspective to model video watch time, from which we propose a white-box statistical framework that directly translates various user behavior assumptions in watching (short) videos into statistical watch time models. These behavior assumptions are portrayed by our domain knowledge on users’ behavior modes in video watching. We further employ bucketization to cope with user’s non-stationary watching probability over the video horizon, which additionally helps to respect the constraint of video length and facilitate the practical compatibility between the continuous regression event of watch time and other binary classification events. We test our models extensively on two public datasets, a large-scale offline industrial dataset, and an online A/B test on a short video platform with hundreds of millions of daily-active users. On all experiments, our models perform competitively against strong relevant baselines, demonstrating the efficacy of our user-centric perspective and proposed framework.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The popularization of online video platforms, such as YouTube Shorts, TikTok, and Instagram Reels, has rendered the rising importance of (short) video recommendation for effective large-scale information filtering and entertainment systems (Zhou et al., 2010; Davidson et al., 2010; Tang et al., 2017; Liu et al., 2019, 2021; Gao et al., 2022b). On these platforms, video watch time emerges as a central feedback signal, which indicates users’ interests and serves as a core optimization metric for companies’ commercialization (Cai et al., 2023a, b). This metric is further highlighted by the relative inappropriateness of classical engagement signals such as click-through rate, since recommended videos on these platforms are typically (re-)played automatically and hence user’s click action is erased. To improve user engagement on the platform, accurate prediction of a user’s watch time on a video has become a central task in the ranking stage, which nevertheless has been challenging (Zhao et al., 2023b; Sun et al., 2024; Bai et al., 2024; Lin et al., 2024).
@@
Unlike conventional binary prediction tasks, such as estimating click-through rate or like rate, video watch time prediction is significantly more complex, making current state-of-the-art methods still less satisfactory. The complexity is not solely due to its basis in continuous signal estimation; rather, it is significantly influenced by the complex user-video interaction during video playback (Zhao et al., 2023a). Concretely, users may exhibit different behavior modes in watching recommended videos. For example, some users may keep “jumping” on the video progress bar. Some others, however, may watch the video sequentially, but quit in interim (or till finishing). Further, users’ interests/retention in the video, formulated as the probability of watching a video second, may be non-stationary and keep changing over the video horizon. Besides, each video has its own length, which naturally constraints video watch time and constitutes a layer of difficulty in modeling watch time (Zhan et al., 2022; Zheng et al., 2022).
@@
Despite the importance and challenges, prior works on modeling video watch time tend to overlook the importance of understanding user behavior in video watching, and typically focus on relatively black-box mechanical enhancements of the classical regression/classification objectives. By its continuous and scalar-valued nature, video watch time is classically modeled as an ordinary regression problem (Zhan et al., 2022; Tang et al., 2023), whose performance is however limited due to data outliers and skewed distribution (Sun et al., 2024). Covington et al. (2016) convert this regression problem into learning the odds of the video click-through-rate in the click-through classification, by using the observed video watch time as the per-sample weight. Zhan et al. (2022) split training data evenly by video duration, and learn a regression model for each duration group for group-wise watch-time quantiles. Lin et al. (2023) break the regression problem of video watch time into a sequence of classification problems arranged in a tree structure, motivated by regression trees (Hastie et al., 2009). Though some of these methods have been applied to industrial systems, they still leave clear room for further improvement, particularly on suiting watch time prediction to user behavior and accounting for the complexity of user-video interaction.
@@
In this paper, we for the first time take on a user-centric perspective to advance video watch time prediction, from which we propose a white-box statistical framework for modeling user’s video watch time. In particular, to factor in users’ various behavior modes, we distill domain knowledge of user behavior into various behavioral assumptions of watching (short) videos. These behavioral assumptions are then directly translated into statistical models of video watch time. To cope with user’s non-stationary watching probability over the video horizon, we bucketize video horizon into buckets/bins and learn an individual watching probability for each bucket. The flexibility of bucketization enables straightforward statistical models without sacrificing overall model capacity. Bucketization can further implicitly satisfy the constraint of video length, since short videos should have watching probabilities concentrated on lower-value buckets (e.g., 5-10s) and long videos on higher-value buckets (e.g., 95-100s). With these considerations, our framework allows a unified principled statistical model on videos of all lengths.
@@ 2.1. Video Watch Time Prediction
Video watch time prediction aims to predict the watch time of a user on a video, given user and video information (Sun et al., 2024). By the scalar nature of watch time, one can intuitively train a model by standard regression loss, which however suffers from data outliers, skewed distribution, and incompatibility with classification-based events (Sun et al., 2024). Covington et al. (2016) improve YouTube video recommendation by using ground-truth video watch time as the per-sample weight to enhance the logistic regression of predicting impressed videos. A watch time prediction can then be conveniently obtained by exponentiating the learned odds, which suits well to the long-tail nature of video-watch time. This approach, however, is only valid under the assumption that the rate of viewed/impressed video is low, and is thus inappropriate to most real-world short video platforms where videos are played automatically and hence all are viewed. Continuing this research thread, TPM (Lin et al., 2023) constructs a tree-like structure to transform the task of video watch time prediction into a series of classification problem, for directing the given sample to its corresponding leaf node. The predicted watch time is then obtained by a simple model within each leaf node, such as the average of training samples within that leaf. Sun et al. (2024) too predict video watch time through multiple classification problems that are constructed by their proposed discretization method for better balancing the learning and restoration errors. Lin et al. (2024) utilize quantile regression (Koenker, 2005) to capture the watch time distribution, for providing more accurate and robust estimates.
@@ 2.2. Classification for Regression
There has been a long history of tackling regression problems from a classification lens, especially when the precise numerics of the target is less important. A classical example of this approach is ordinal regression, which models the ordinal nature of the data (McCullagh, 1980) and is typically extended from classical classification models, e.g., logistic regression (McCullagh, 1980), SVM (Shashua and Levin, 2002), and decision tree (Frank and Hall, 2001). Ordinal regression has been applied to various applications, including age estimation (Niu et al., 2016; Beckham and Pal, 2017), depth estimation (Fu et al., 2018; Diaz and Marathe, 2019), and pose estimation (Hsu et al., 2018). Classification has also been used to model integer-valued data, where each distinct value is treated as a different class (Waugh, 1995). In computer vision, modeling pixels as discrete valued via classification rather than continuous valued has been well adopted for its representational and training advantages (Van Den Oord et al., 2016, 2017; Oord et al., 2018)
@@ 3. The Proposed Method: SWaT
As discussed in Section 1, we statistically model video watch time based on our distinct user-centric perspective. Concretely, to suit users’ different behavior modes in (short) video watching, we build our statistical models by directly translating different user behavior assumptions, which are a distillation of our domain knowledge. Further, considering the users’ varying interests/retention intention over the video horizon, we bucketize the infinite video horizon onto smaller buckets, which allows intuitive statistical modeling on bucket-specific watching probabilities while ensures model capacity and implicitly respects the video length constraint. In inference time, our models permit efficient closed-form point estimates of video watch time via the expectation of the specified statistical models. Additional benefits of bucketization strategy on stabilizing training objective is discussed in details in Section 3.4.
@@ 3.1. Action Steps of SWaT
The first action step in our SWaT framework is to partition infinite video horizon $[0,\infty)$ by $N$ endpoints $\left\{x_{i}\right\}_{i=1}^{N}$, which facilitates specifying user behavior assumptions and learning non-stationary watching probabilities. Notationally, we have: $-\epsilon=x_{0}<x_{1}<x_{2}<\cdots<x_{N}<x_{N+1}=\infty$, where $-\epsilon\approx 0$ to conveniently include $t=0$ into the first bucket. The $i^{\mathrm{th}}$ bucket is denoted as $B_{i}$ and the width of $B_{i}$ is $\Delta_{i}$, i.e., $B_{i}=(x_{i-1},x_{i}],\Delta_{i}=x_{i}-x_{i-1},\forall\,i=1,\ldots,N$.
With bucketization, different user behavior is factored in our statistical models by different behavioral assumptions that portrays users’ different behavior modes of watching (short) videos within each bucket. These behavioral assumptions are then mapped to statistical models for video watch time, which depend on bucket-specific watching probability $p_{i}$ for each bucket $B_{i}$. Value variation of learned $\left\{\widehat{p}_{i}\right\}$ demonstrates user’s non-stationary interest/probability of continuing watching the video over its horizon.
With the observed video watch time $t$, our statistical models, in particular, $\{p_{i}\}$, are fitted by maximum likelihood estimation (MLE, Ferguson, 1996; Casella and Berger, 2001). After fitting the model, in inference time, we estimate the expected watch time of a user on a video in closed-form based on the specified statistical model and the learned $\left\{\widehat{p}_{i}\right\}$, typically by summing up the user’s predicted watch time in each bucket $B_{i}$, i.e., our estimate $\widehat{\mu}[T]$ of expected video watch time $\mathbb{E}[T]$ is
@@ 3.2. SWaT-Binom: Binomial Model for Random Pick-and-Play
Under Assumption 1, in particular, identical watching probability in $B_{i}$ and independence between the selection of $i^{\mathrm{th}}$ second and $(i+1)^{\mathrm{th}}$ second, the watch time in each bucket $B_{i}$ can be modeled as a binomial distribution. A pictorial illustration is in Fig. 2. Technically, to avoid undesirable value truncation, $x_{N}$ should be chosen sufficiently large to cover the range of $T$. The last bucket $[x_{N},x_{N+1}=\infty)$ is not used due to its infinite support. For an observation $t$ of $T$, denote the user’s watch time in $B_{i}$ as $t_{i}$, where $\sum_{i=1}^{N}t_{i}=t$. We use MLE to fit the watching probability $p_{i}$. Concretely, for each $B_{i}$, with the given $\Delta_{i}$ and $t_{i}$, we have
@@
For the construction of (soft) label $t_{i}/\Delta_{i}$, since user’s real per-bucket watch time $\left\{t_{i}\right\}_{i=1}^{N}$ is typically not recorded in current industrial recommendation systems. We adopt an minimalist assumption that the recorded users view the video stream sequentially, except in the last watched bucket, where the user gets bored and only picks some seconds to watch, and afterwards stops watching. We note that this “boredom” assumption is not exerted in inference time, and is a practical compromise that can be removed by improved system logging. With this extra assumption, we use the following construction of the label $t_{i}/\Delta_{i}$ based on the observation $t$: if $t\in B_{i}\iff x_{i-1}<t\leq x_{i}$, we have $t_{j}=\Delta_{j},\forall\,j=1,\ldots,i-1$ and $t_{j}=0,\forall\,j>i$. The (soft) labels $\left\{t_{i}/\Delta_{i}\right\}_{i=1}^{N}$ are then constructed as,
@@
Eq. 2 is also used as piecewise linear encoding of integer value feature for tabular deep learning (Gorishniy et al., 2022), which has the benefit of allowing exact recovery of the feature value rather than the nearest bucket endpoint.
@@
Industrial recommendation systems are typically jointly trained by many events that are mostly classification-based, e.g., like and share. ${\mathcal{L}}_{\text{\sc SWaT-Binom}}$ transforms the regression event of watch time into (soft) classification, facilitating its compatibility with others.
@@ 3.3.1. Non-stationary Watching Probabilities over Video Horizon
Under Assumption 2, user’s video watch time can be modeled as a bucketized geometric distribution. Specifically, for a given set of watching probabilities $\left\{p_{i}\right\}_{i=1}^{N+1}$, if the observation $t\in B_{i}=(x_{i-1},x_{i}]$, the probability of $T=t$ is
(4) $\Pr\left(T=t\right)=p_{i}^{t-x_{i-1}}\times(1-p_{i})\times\prod_{j=1}^{i-1}p_{j}^{\Delta_{j}}\,,$
@@
$\quad\log\left(1-p_{i}\right)+\sum_{j=1}^{i-1}\Delta_{j}\log p_{j}\,,$
where the index $i$ corresponds to the bucket $B_{i}$ that the given training sample falls into and may differ across samples.

### v2

@@ Abstract.
The significance of estimating video watch time has been highlighted by the rising importance of (short) video recommendation, which has become a core product of mainstream social media platforms. Modeling video watch time, however, has been challenged by the complexity of user-video interaction, such as different user behavior modes in watching the recommended videos and varying watching probability over the video progress bar. Despite the importance and challenges, existing literature on modeling video watch time mostly focuses on relatively black-box mechanical enhancement of the classical regression/classification losses, without factoring in user behavior in a principled manner. In this paper, we for the first time take on a user-centric perspective to model video watch time, from which we propose a white-box statistical framework that directly translates various user behavior assumptions in watching (short) videos into statistical watch time models. These behavior assumptions are portrayed by our domain knowledge on users’ behavior modes in video watching. We further employ bucketization to cope with user’s non-stationary watching probability over the video progress bar, which additionally helps to respect the constraint of video length and facilitate the practical compatibility between the continuous regression event of watch time and other binary classification events. We test our models extensively on two public datasets, a large-scale offline industrial dataset, and an online A/B test on a short video platform with hundreds of millions of daily-active users. On all experiments, our models perform competitively against strong relevant baselines, demonstrating the efficacy of our user-centric perspective and proposed framework.
Video Watch Time Prediction; Short Video Recommendation; User Behavior Analysis
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709415isbn: 979-8-4007-1245-6/25/08ccs: Information systems Collaborative and social computing systems and toolsccs: Information systems Data miningccs: Human-centered computing User models
@@ 1. Introduction
The popularization of online video platforms, such as YouTube Shorts, TikTok, and Instagram Reels, has rendered the rising importance of (short) video recommendation as effective large-scale information filtering and entertainment systems (Zhou et al., 2010; Davidson et al., 2010; Tang et al., 2017; Liu et al., 2019, 2021; Gao et al., 2022b). On these platforms, video watch time emerges as a central feedback signal, which indicates users’ interests and serves as a core optimization metric for companies’ commercialization (Cai et al., 2023a, b). This metric is further highlighted by the relative inappropriateness of classical engagement signals such as click-through rate, since recommended videos on these platforms are typically (re-)played automatically and hence user’s click action is erased. To improve user engagement on the platform, accurate prediction of a user’s watch time on a video has become a central task in the ranking stage, which nevertheless has been challenging (Zhao et al., 2023b; Sun et al., 2024; Bai et al., 2024; Lin et al., 2024).
@@
Unlike conventional binary prediction tasks, such as estimating click-through rate or like rate, video watch time prediction is significantly more complex, making current state-of-the-art methods still less satisfactory. The complexity is not solely due to its basis in continuous signal estimation; rather, it is significantly influenced by the complex user-video interaction during video playback (Zhao et al., 2023a). Concretely, users may exhibit different behavior modes in watching recommended videos. For example, some users may keep “jumping” on the video progress bar. Some others, however, may watch the video sequentially, but quit in interim (or till finishing). Further, users’ interests/retention in the video, formulated as the probability of watching a video second, may be non-stationary and keep changing over the video progress bar. Besides, each video has its own length, which naturally constraints video watch time and constitutes a layer of difficulty in modeling watch time (Zhan et al., 2022; Zheng et al., 2022).
@@
Despite the importance and challenges, prior works on modeling video watch time tend to overlook the importance of understanding user behavior in video watching, and typically focus on relatively black-box mechanical enhancements of the classical regression and/or classification objectives. By its continuous and scalar-valued nature, video watch time is classically modeled as an ordinary regression problem (Zhan et al., 2022; Tang et al., 2023), whose performance is however limited due to data outliers and skewed distribution (Sun et al., 2024). Covington et al. (2016) convert this regression problem into learning the odds of the video click-through-rate in the click-through classification, by using the observed video watch time as the per-sample weight. Zhan et al. (2022) split training data evenly by video duration, and learn a regression model for each duration group for group-wise watch-time quantiles. Lin et al. (2023) break the regression problem of video watch time into a sequence of classification problems arranged in a tree structure, motivated by regression trees (Hastie et al., 2009). Though some of these methods have been applied to industrial systems, they still leave clear room for further improvement, particularly on suiting watch time prediction to user behavior and accounting for the complexity of user-video interaction.
@@
In this paper, we for the first time take on a user-centric perspective to advance video watch time prediction, from which we propose a white-box statistical framework for modeling user’s video watch time. In particular, to factor in users’ various behavior modes, we distill domain knowledge of user behavior into various behavioral assumptions of watching (short) videos. These behavioral assumptions are then directly translated into statistical models of video watch time. To cope with user’s non-stationary watching probability over the video progress bar, we bucketize video progress bar into buckets/bins and learn an individual watching probability for each bucket. The flexibility of bucketization enables straightforward statistical models without sacrificing overall model capacity. Bucketization can further pad videos to a unified maximum length but implicitly respect the constraint of each video’s length, since short videos could have watching probabilities concentrated on lower-value buckets (e.g., 5-10s) and long videos on higher-value buckets (e.g., 95-100s). With these considerations, our framework allows a unified principled statistical model on videos of all lengths.
@@ 2.1. Video Watch Time Prediction
Video watch time prediction aims to predict a user’s watch time on a video, given user and video information (Sun et al., 2024). By the scalar nature of watch time, one can intuitively train a model by standard regression loss, which however suffers from data outliers, skewed distribution, and incompatibility with classification-based events (Sun et al., 2024). Covington et al. (2016) improve YouTube video recommendation by using ground-truth video watch time as the per-sample weight to enhance the logistic regression of predicting impressed videos. A watch time prediction can then be conveniently obtained by exponentiating the learned odds, which suits well to the long-tail nature of video-watch time. This approach, however, is only valid under the assumption that the rate of viewed/impressed video is low, and is thus inappropriate to most real-world short video platforms where videos are played automatically and hence all are viewed. Continuing this research thread, TPM (Lin et al., 2023) constructs a tree-like structure to transform the task of video watch time prediction into a series of classification problems, for directing the given sample to its corresponding leaf node. The predicted watch time is then obtained by a simple model within each leaf node, such as the average of training samples within that leaf. Sun et al. (2024) too predict video watch time through multiple classification problems that are constructed by their proposed discretization method for better balancing the learning and restoration errors. Lin et al. (2024) utilize quantile regression (Koenker, 2005) to capture the watch time distribution, for providing more accurate and robust estimates.
@@ 2.2. Classification for Regression
There has been a long history of tackling regression problems from a classification lens, especially when the precise numerics of the target is less important. A classical example of this approach is ordinal regression, which models the ordinal nature of the data (McCullagh, 1980) and is typically extended from classical classification models, e.g., logistic regression (McCullagh, 1980), SVM (Shashua and Levin, 2002), and decision tree (Frank and Hall, 2001). Ordinal regression has been applied to various applications, including age estimation (Niu et al., 2016; Beckham and Pal, 2017), depth estimation (Fu et al., 2018; Diaz and Marathe, 2019), and pose estimation (Hsu et al., 2018). Classification has also been used to model integer-valued data, where each distinct integer is treated as a different class (Waugh, 1995). In computer vision, modeling pixels as discrete values via classification rather than continuous values has been well adopted for its representational and training advantages (Van Den Oord et al., 2016, 2017; Oord et al., 2018).
@@ 3. The Proposed Method: SWaT
As discussed in Section 1, we statistically model video watch time based on our distinct user-centric perspective. Concretely, to suit users’ different behavior modes in watching (short) videos, we build our statistical models by directly translating different user behavior assumptions, which are distillation of our domain knowledge. To relieve the modeling burden while allow aligning with the video length constraint, our statistical models operate on the entire video horizon: the (infinitely long) video progress bar counting in possibly indefinite time automatic replays, which conceptually pads out all possible time seconds in a video. Further, considering the users’ varying interests/retention intention over the video horizon, we bucketize the infinite video horizon onto smaller buckets, which allows intuitive statistical modeling on bucket-specific watching probabilities while ensures model capacity and the realistic finite-time video replays. In inference time, our models permit efficient closed-form point estimates of video watch time via the expectation of the specified statistical models. Additional benefits of the bucketization strategy on stabilizing training objective is discussed in details in Section 3.4.
@@ 3.1. Action Steps of SWaT
The first action step in our SWaT framework is to partition the infinite video horizon $[0,\infty)$ by $N$ endpoints $\left\{x_{i}\right\}_{i=1}^{N}$, which facilitates specifying user behavior assumptions and learning non-stationary watching probabilities. Notationally, we have: $-\epsilon=x_{0}<x_{1}<x_{2}<\cdots<x_{N}<x_{N+1}=\infty$, where $-\epsilon\approx 0$ to conveniently include $t=0$ into the first bucket. The $i^{\mathrm{th}}$ bucket is denoted as $B_{i}$ and the width of $B_{i}$ is $\Delta_{i}$, i.e., $B_{i}=(x_{i-1},x_{i}],\Delta_{i}=x_{i}-x_{i-1},\forall\,i=1,\ldots,N$.
With bucketization, different user behavior is factored in our statistical models by different behavioral assumptions that portray users’ different behavior modes of watching (short) videos within each bucket. These behavioral assumptions are then mapped to statistical models for video watch time, which depend on bucket-specific watching probability $p_{i}$ of each bucket $B_{i}$. Value variation of the learned $\left\{\widehat{p}_{i}\right\}$ manifests user’s non-stationary interest/probability of continuing watching the video over its horizon.
With the observed video watch time $t$, our statistical models, in particular, $\{p_{i}\}$, are fitted by maximum likelihood estimation (MLE, Ferguson, 1996; Casella and Berger, 2001). After fitting the model, in inference time, we estimate the expected watch time of a user on a video in closed-form based on the specified statistical model and the learned $\left\{\widehat{p}_{i}\right\}$, typically by summing up the user’s predicted watch time in each bucket $B_{i}$, i.e., our estimate $\widehat{\mu}[T]$ of the expected video watch time $\mathbb{E}[T]$ is
@@ 3.2. SWaT-Binom: Binomial Model for Random Pick-and-Play
Under Assumption 1, in particular, identical watching probability in $B_{i}$ and independence between watching the $i^{\mathrm{th}}$ second and the $(i+1)^{\mathrm{th}}$ second, the watch time in each bucket $B_{i}$ can be modeled as a binomial distribution. A pictorial illustration is in Fig. 2. Technically, to avoid undesirable value truncation, $x_{N}$ should be chosen sufficiently large to cover the range of $T$ in practice. The last bucket $[x_{N},x_{N+1}=\infty)$ is not used due to its infinite support. For an observation $t$ of $T$, denote the user’s watch time in $B_{i}$ as $t_{i}$, where $\sum_{i=1}^{N}t_{i}=t$. We use MLE to fit the watching probability $p_{i}$. Concretely, for each $B_{i}$, with the given $\Delta_{i}$ and $t_{i}$, we have
@@
For the construction of (soft) label $t_{i}/\Delta_{i}$, since user’s real per-bucket watch time $\left\{t_{i}\right\}_{i=1}^{N}$ is typically not recorded in current industrial recommendation systems. We adopt an minimalist assumption that the recorded users view the video stream sequentially, except in the last watched bucket, where the user gets bored and only picks some seconds to watch, and afterwards stops watching. We note that this “boredom” assumption is not exerted in inference time, and is a practical compromise that can be removed by improved system logging. With this extra assumption, we use the following construction of the label $t_{i}/\Delta_{i}$ based on the observation $t$: if $t\in B_{n}\iff x_{n-1}<t\leq x_{n}$, we have $t_{i}=\Delta_{i},\forall\,i=1,\ldots,n-1$ and $t_{i}=0,\forall\,i>n$. The (soft) labels $\left\{t_{i}/\Delta_{i}\right\}_{i=1}^{N}$ are then constructed as,
@@
Eq. 2 is also used as piecewise linear encoding of an integer value feature for tabular deep learning (Gorishniy et al., 2022), which has the benefit of allowing exact recovery of the feature value rather than the nearest bucket endpoint.
@@
Industrial recommendation systems are typically jointly trained by many events that are mostly classification-based, e.g., like and share. ${\mathcal{L}}_{\text{\sc SWaT-Binom}}$ transforms the regression event of video watch time into (soft) classification, facilitating its compatibility with others in model training.
@@ 3.3.1. Non-stationary Watching Probabilities over Video Horizon
Under Assumption 2, user’s video watch time can be modeled as a bucketized geometric distribution. Specifically, for a given set of watching probabilities $\left\{p_{i}\right\}_{i=1}^{N+1}$, if the observation $t\in B_{n}=(x_{n-1},x_{n}]$, the probability of $T=t$ is
(4) $\Pr\left(T=t\right)=p_{n}^{t-x_{n-1}}\times(1-p_{n})\times\prod_{i=1}^{n-1}p_{i}^{\Delta_{i}}\,,$
@@
$\quad\log\left(1-p_{n}\right)+\sum_{i=1}^{n-1}\Delta_{i}\log p_{i}\,,$
where the index $n$ refers to the bucket $B_{n}$ that the given training sample falls into and may differ across samples.